# Structural Alignment Improves Graph Test-Time Adaptation

### v1

@@ D.1 Datasets Setup
The synthetic dataset is generated by the contextual stochastic block model (CSBM) (Deshpande et al., 2018). The generated graph contains 6000 nodes and 3 classes. We alter the edge connection probability matrix $\mathbf{B}$ and the label ratio $\mathbb{P}_{Y}$ Specifically, the edge connection probability matrix $\mathbf{B}$ is a symmetric matrix given by
@@
Under the study of different PU levels we use the data from signal $gg$. Compared to the Pileup datasets used in the (Liu et al., 2024), we explicitly label all types of particles to better align the neighborhood distribution. Note that the charged particle can achieve all most perfect recall as their label information is encoded in their feature, hence the four-class classification still reduces into a binary classification between LC neutral particles and OC neutral particles. For the study of different data generation process, we use PU10 from signal $gg\rightarrow qq$ and $qq\rightarrow gg$.
D.2 Pretraining Setup
Model architecture and Pretraining
We use GraphSAGE as the model backbone, with a 3-layer mean pooling GNN as the encoder and a 2-layer MLP with a batch normalization layer as the classifier. The hidden dimension of the encoder is set to 300 for Arxiv and MAG, 50 for Pileup, 128 for DBLP/ACM, and 20 for the synthetic datasets. The hidden dimension of the classifier is set to 300 for Arxiv and MAG, 50 for Pileup, 40 for DBLP/ACM, and 20 for the synthetic datasets. For the GPRGNN backbone, we follow the same configuration but increase the number of encoder layers to 5. All models are trained for 400 epochs with the learning rate set to 0.003. A learning rate scheduler is applied with a decay rate of 0.9. All experiments are repeated 5 times under different initializations and data splits to ensure consistency.
D.3 Evaluation and Metric
The source graph is splitted into train/val/test 60/20/20 during the pretraining stage. The target graph is splitted into labeled/unlabeled 3/97 during test time. We use the the labeled nodes in the target graph to do hyperparameter tuning and select the model with the optimal validation score. We follow the metrics in (Liu et al., 2024) for evaluation. MAG, Arxiv, DBLP/ACM, and synthetic datasets are evaluated with accuracy. For the MAG datasets, only the top 19 classes are evaluated. Pileup is evaluated with f1 score.
D.4 Hyperparameter Tuning
@@
  • AdaRC introduces two hyperparameters learning rate $lr$ and the number of epochs $T$. Based on their hyperparameter study, We select $lr$ from $\{1,5,10\}$ and $T$ from $\{1,10,100\}$.
@@
  • SOGA introduces one hyperparameters $lr$ for one epoch update. We select $lr$ from $\{0.001,0.01\}$.
  • TENT introduces one hyperparameters $lr$ for one epoch update. We select $lr$ from $\{0.001,0.01,0.05\}$.
  • T3A introduces one hyperparameters $M$ for deciding the number of supports to restore. We select $M$ from $\{5,20,50,100\}$

### v2

@@ F.1 Datasets Setup
The synthetic dataset is generated by the contextual stochastic block model (CSBM) [60]. The generated graph contains 6000 nodes and 3 classes. We alter the edge connection probability matrix $\mathbf{B}$ and the label ratio $\mathbb{P}_{Y}$ Specifically, the edge connection probability matrix $\mathbf{B}$ is a symmetric matrix given by
@@
Under the study of different PU levels we use the data from signal $gg$. Compared to the Pileup datasets used in the [59], we explicitly label all types of particles to better align the neighborhood distribution. Note that the charged particle can achieve all most perfect recall as their label information is encoded in their feature, hence the four-class classification still reduces into a binary classification between LC neutral particles and OC neutral particles. For the study of different data generation process, we use PU10 from signal $gg\rightarrow qq$ and $qq\rightarrow gg$.
F.2 Pretraining Setup
Model Architecture and Pretraining.
We use GraphSAGE as the model backbone, with a 3-layer mean pooling GNN as the encoder and a 2-layer MLP with a batch normalization layer as the classifier. The hidden dimension of the encoder is set to 300 for Arxiv and MAG, 50 for Pileup, 128 for DBLP/ACM, and 20 for the synthetic datasets. The hidden dimension of the classifier is set to 300 for Arxiv and MAG, 50 for Pileup, 40 for DBLP/ACM, and 20 for the synthetic datasets. For the GPRGNN backbone, we follow the same configuration but increase the number of encoder layers to 5. For the GCN backbone, we use a 3-layer GCN as the feature encoder, followed by the same 2-layer MLP classifier as in the other backbones. All models are trained for 400 epochs with the learning rate set to 0.003. We use Adam [75] as the optimizer for all models, and apply a learning rate scheduler with a decay rate of 0.9. All experiments are repeated 5 times under different initializations and data splits to ensure consistency.
Hardware and Frameworks.
All experiments are conducted on a server with a single NVIDIA RTX 6000 GPU and an AMD EPYC 7763 64-Core Processor. We use PyTorch 2.2.2 [76] with CUDA 12.1 as the framework. For graph neural networks, we use PyTorch Geometric 2.5.2 [77].
F.3 Evaluation and Metric
The source graph is splitted into train/val/test 60/20/20 during the pretraining stage. The target graph is splitted into labeled/unlabeled 3/97 during test time. We use the the labeled nodes in the target graph to do hyperparameter tuning and select the model with the optimal validation score. We follow the metrics in [59] for evaluation. MAG, Arxiv, DBLP/ACM, and synthetic datasets are evaluated with accuracy. For the MAG datasets, only the top 19 classes are evaluated. For the Pileup datasets, we evaluate the performance with f1 score.
F.4 Hyperparameter Tuning
@@
  • Matcha introduces two hyperparameters learning rate $lr$ and the number of epochs $T$. Based on their hyperparameter study, We select $lr$ from $\{0.1,1,5,10\}$ and $T$ from $\{2,10\}$.
@@
  • HomoTTT introduces two hyperparameter $lr$ and $k0$. The hyperparameter $k0$ is a scaling factor for the model selection score. We present the results updated by one epoch. We select $lr$ from $\{0.001,0.01\}$ and $k0$ from $\{0.001,0.01\}$.
  • SOGA introduces one hyperparameter $lr$. We present the results updated by one epoch and select $lr$ from $\{0.001,0.01\}$.
  • TENT introduces one hyperparameter $lr$. We present the results updated by one epoch and select $lr$ from $\{0.001,0.01,0.05\}$.
  • T3A introduces one hyperparameter $M$ for deciding the number of supports to restore. We select $M$ from $\{5,20,50,100\}$
  • ActMAD introduces one hyperparameter $lr$. We present the results updated by one epoch and select $lr$ from $\{0.001,0.01\}$. We adapt ActMAD to GNNs by aligning the outputs of each GNN layer before the classifier.